# Extra dimensional axion patterns

### v1

@@ Abstract
We study for the first time the complete parameter space of a bulk extra-dimensional axion. We find novel regimes where no single KK mode is produced along the canonical QCD axion line, and instead, it is maximally deviated along with several other axions that constitute a multiple solution to the strong CP problem. In the most common extra-dimensional models, namely for flat and curved Randall-Sundrum scenarios, and assuming that all Peccei-Quinn breaking comes from QCD, we find that these solutions are however subject to tight phenomenological constraints. In light of these results, only one – canonical – pattern can be expected from a bulk axion in one or more extra spacetime dimensions. As a byproduct, we generalize the axions eigenvalue and eigenvector equations for an arbitrary number of spacetime dimensions and compactifications.
@@ I Motivation and summary
Phenomenological studies of a 5D QCD axion have been presented in a series of seminal works [8, 9, 10] and have garnered renewed attention in recent years [11, 12, 13, 14, 15, 16, 17]. One motivation for this framework is to produce a weak axion coupling to gauge bosons from a high fundamental PQ scale, by the same mechanism that lowers the Planck scale in the higher dimensional theory. Importantly, this framework also allows to decouple the mass of the zero mode from the PQ scale, $m_{\rm PQ}$, which can be instead set by the scale of the extra dimension, $R^{-1}$, as first pointed out in Ref. [8]. In this work, we explore novel consequences of this large $R$ limit and the critical role of the massive KK modes in solving the strong CP problem.
@@
We will prove that an extra-dimensional bulk axion can either give rise to $1$ canonical axion or a set of $N$ deviated QCD maxions predicted by the sum rule (I.1), that is maxions are the non-canonical solutions of generic extra spacetime dimensional models (assuming that QCD is the only source of PQ breaking). This conclusion holds for any number of dimensions, in either flat [18, 19] or curved Randall-Sundrum (RS) [20, 21] scenarios. The two possible patterns that can arise are shown schematically in Fig. 1, represented by an orange or a blue line. The first indicates a canonical QCD axion solution, where all but one mode are decoupled from the solution to the strong CP problem. The second type of pattern – discussed, to the best of our knowledge, for the first time in this work – corresponds to the extra-dimensional QCD maxion scenario, where no single axion is found close to the canonical band. As we will show, such patterns arise commonly in the regime $R^{-1}\lesssim m_{\rm PQ}$, and encode a mapping to the properties of the fundamental higher-dimensional bulk axion.
Some of these patterns could be generated by other frameworks, namely the string axiverse [22] and Grand Unified Theories [23], possibly in regimes not yet explored. Our ultimate goal would be to find which exotic patterns survive the scrutiny of phenomenological probes, that constrain differently the UV models, in order to finally point towards new regions of signal for the QCD axion(s) and their origin. In this work, we take the first step in this direction, by exploring the variety of constraints on exotic multiple axion patterns from models with extra spacetime dimensions.
@@
The main conclusion of our work is that the QCD maxion patterns from a PQ bulk field (in generic extra-dimensional models) are tightly constrained by a combination of consistency, astrophysical and gravitational bounds. This holds independently of the number of (universal) spacetime dimensions and the VEV profile of the PQ field in the bulk. We, therefore, expect that more complex constructions, with additional bulk fields or less generic compactifications, are required to produce the most exotic patterns in Fig. 1. In turn, the observation of a plateau of heavy fields together with a canonical QCD axion could be explained by a single PQ field, propagating in the bulk of a hidden dimension. We will demonstrate how a potential observation of such a plateau could in this case hint towards the location of the zero mode in the canonical QCD axion band.
@@ II Extra dimensional models
with $y_{n}$ denoting a coordinate of the extra dimensions, that we assume to be compactified on a $(S^{1}/\mathbb{Z}^{2})$ orbifold of universal radius $R$. Accordingly, $y_{n}\to-y_{n}$ are identified for all points in the interval $[-\pi R,\pi R]$. In the equation above, $g$ is the determinant of the metric, $\mathcal{R}$ the Ricci scalar, $M_{d}$ the $d$-dimensional Planck mass, and $\Lambda_{B}$ the vacuum energy of the bulk  [^1] We will denote coordinates belonging to the full extra-dimensional models, to the 4D EFT and to the extra-dimensions with capital indices (e.g. $x^{A}$), greek letters (e.g. $x^{\mu}$ with $\mu=0,1,2,3$) and in latin lower case (e.g. $x^{a}$ with $a=5,6,\dots,\,d$, or $y_{n}$ with $n=1,\dots\,,\delta$), respectively. We will also make use of $\mathbf{y}=(y_{1},\dots\,,y_{\delta})$ to highlight the full set of extra-dimensional coordinates.. Regarding the matter content, we focus on models where the SM fields are localized on the branes, that is the boundaries of the extra dimensions.
@@
where the $\delta$-function localizes the axion couplings to QCD on the SM brane, at $y_{n}=\pi R$. The interactions above are assumed to be generated at a fundamental scale $M_{s}$, with $f_{d}$ denoting the PQ breaking scale. The former can be removed from the kinetic term by an appropriate redefinition of the axion field and the axion decay constant.
@@ III General mass matrix
$m_{\rm PQ}^{2}\equiv\chi_{\rm QCD}/f_{\rm PQ}^{2}$ is the standard PQ mass and $\chi_{\text{QCD}}$ is the QCD topological susceptibility. Such mass matrix reproduces any of the models described in the previous section, by simply matching the WFs to Eqs. (II.17) (flat) and (II.28) (RS). Note that the main difference between these constructions in the $\mu\gg 1$ limit lies not in the mass ratios, but rather in the expressions for these WFs, which are exponentially enhanced by the curvature.
@@
${\sum\limits_{n=0}^{\infty}\dfrac{(\psi_{n})^{2}}{\lambda^{2}-(\mu_{n}/\mu_{1})^{2}\mathsf{y}^{2}}=1\,;}$ (III.4)
see App. A for details. A more general approach will also be derived in Sec. VI.
@@ IV Maximal displacements of KK axions
The $g$-factors can now be found using Eqs (IV.4) and (III.4):
@@
Remarkably, using the definition above, one rediscovers the QCD axion sum rule
@@
proven in all generality in Ref. [5] (see App. B for details).
We will now infer the consequences of these general results to the models introduced in the previous section.
@@ IV.1 QCD maxions in flat space
In flat scenarios, the sum in Eq. (III.4) can be performed exactly:
@@
These results confirm our expectations.
@@ IV.2 QCD maxions in curved space
In the RS model, the computations are more elaborate as $\mu_{n}\propto\gamma_{n}$. The sum in Eq. (III.4) can be performed exactly by means of the identity [24]
@@
These results, together with Eq. (IV.33), allow us to estimate the localization of the lightest eigenvalue, as well as of the plateau, for the several scenarios represented in Fig. 2(b).
@@ V Theoretical and experimental constraints
Figure 3: Collection of bounds on $\psi_{1}/\psi_{0}$ as a function of the lightest graviton mass, $\mu_{1}$. The red, green and purple regions are excluded by fifth force experiments [27, 28, 29, 30, 31, 32, 33, 34], astrophysics [35, 36] and collider searches [25, 37, 38], respectively. The black dashed line corresponds to the threshold mass value below which KK maxions can be generated; see Eq. (V.8). This assumes that the KK tower remains perturbative up to TeV energies.
To combine these constraints, motivated by the form of the eigenvalue equation (III.4), we consider a simplified approach to extra-dimensional maxions, that relies on using the $2\times 2$ block mass matrix
@@
This is equivalent to truncating the sum in Eq. (III.4) to the second term. The lightest eigenvalue can then be predicted up to an accuracy of $\mathcal{O}(\mu_{1}/\mu_{2})^{2}$ and, in the limit $\mathsf{y},\psi_{0}\ll\psi_{1}$, reads:
@@
we obtain $g_{a\gamma\gamma\,,\text{plateau}}\approx\,{2.5\times 10^{-10}}\,{\text{GeV}}^{-1}$, after fixing $\mu_{1}\sim 10^{-2}$ eV, $\Lambda\sim 1$ TeV and $|C_{a\gamma}|\approx 1.92$ since all axions inherit the same model-independent coupling to photons from the original interaction in Eq. (II.4). Such coupling is on the verge of the constraints imposed by single-axion searches in stars; the production of additional eigenstates is expected to strengthen this limit severely. Note that model-dependent interactions could be introduced in the IR brane to make the KK axions more photophobic, at the cost of some tuning (see e.g. Ref. [39]). However, supernova and Planck data [40] exclude already the gluonic coupling in Eq. (V.10) for the largest possible displacements identified in Eq. (V.9). Larger values of $\mu_{1}$, which could be realized in the RS framework, would lead to even stronger couplings of the KK axions in the plateau.
@@
We can therefore conclude that the maxion patterns from the extra-dimensional models analysed in this work are in severe tension with the joint combination of gravity and axion constraints.

### v2

# Extra-dimensional axion patterns

@@ Abstract
We study the complete parameter space of a bulk axion in flat and warped extra spacetime dimensions. We characterize in detail the regimes where no single KK mode is produced along the canonical QCD axion line, and instead, it is maximally deviated along with several other axions that constitute a multiple solution to the strong CP problem. In both flat and Randall-Sundrum scenarios, and assuming that all Peccei-Quinn breaking comes from QCD, we find that these solutions are however subject to tight phenomenological constraints. In light of these results, we expect that only KK canonical patterns (with the zero-mode close to the standard QCD line) can emerge from a bulk axion in one or more extra spacetime dimensions. As a byproduct, we generalize the axions eigenvalue and eigenvector equations for an arbitrary number of spacetime dimensions and compactifications.
@@ I Motivation and summary
Phenomenological studies of a 5D QCD axion have been presented in a series of seminal works [8, 9, 10] and have garnered renewed attention in recent years [11, 12, 13, 14, 15, 16, 17]. One motivation for this framework is to produce a weak axion coupling to gauge bosons from a high fundamental PQ scale, by the same mechanism that lowers the Planck scale in the higher dimensional theory. Importantly, this framework also allows to decouple the mass of the zero mode from the PQ scale, $m_{\rm PQ}$, which can be instead set by the scale of the extra dimension, $R^{-1}$, as first pointed out in Ref. [8]. Moreover, the presence of the KK tower can impact substantially the experimental bounds, as well as provide a mechanism for dynamical dark matter in which the relic abundance is shared among several of the KK fields. These aspects have been discussed extensively in the literature in the context of flat extra-dimensional models [18, 19, 20].
@@
In this work, we will focus on the critical role of the KK tower in solving the strong CP problem in the full parameter space of both flat and curved Randall-Sundrum (RS) [21, 22] scenarios. We will prove that an extra-dimensional bulk axion can either give rise to $1$ canonical axion or a set of $N$ deviated QCD maxions predicted by the sum rule (I.1), that is maxions are the non-canonical solutions of generic extra spacetime dimensional models (assuming that QCD is the only source of PQ breaking). The two possible patterns that can arise are shown schematically in Fig. 1, represented by an orange or a blue line. The first indicates a canonical QCD axion solution, where all but one mode are decoupled from the solution to the strong CP problem. The second type of pattern corresponds to the extra-dimensional QCD maxion scenario, where no single axion is found close to the canonical band. Such maximal mixing regimes arise if $A\,\mu_{1}/m_{\rm PQ}\lesssim 1$, with $\mu_{1}$ denoting the lightest graviton mass and $A$ the warping factor associated with the curvature of the extra dimension. Such curvature can take the QCD maxions further out of the canonical line, and separate significantly the zero-mode from the rest of the KK tower; this is representated by the dashed blue line in Fig. 1.
The aim of our work is to, first, characterize fully the axion patterns stemming from extra dimensions and show how the location of the KK modes encodes a mapping to the fundamental properties of the higher-dimensional bulk axion. In flat extra-dimensional scenarios, the mass spectrum and the KK couplings have been previously computed in a series of works [8, 18, 19]. Nonetheless, we will obtain a general formalism from which we will re-derive these results and extend them to the case of a warped extra dimension. Moreover, we will obtain generalized expressions for the eigenvalues and eigenvector equations that apply to a broader class of higher-dimensional axion models with arbitrary compactifications and number of spacetime dimensions, whose relevance extends beyond the scope of our work.
Secondly, we will take into account the most constraining phenomenological probes, to identify which of these patterns are allowed by current data. We will be mostly interested in understanding whether the extra-dimensional maxion scenarios, with $g_{i}\gg 1$, survive this scrutiny. Such scenarios are the most interesting ones in what concerns the possibility of identifying new regions of signal to search for QCD axions. Moreover, some of the patterns represented in Fig. 1 are common to other frameworks, namely the string axiverse [23] and Grand Unified Theories [24]. (This is not surprising, as these patterns follow directly from the generic QCD axion sum rule.) For this reason, investigating the constraints on KK signals from extra-dimensional models is essential to eventually discriminate the origin of a potential multiple axion pattern observed in Nature.
@@
We will show that the QCD maxion patterns from a PQ bulk field are tightly constrained by a combination of consistency, astrophysical and gravitational bounds. To the best of our knowledge, this result is completely new for scenarios with a warped extra dimension. Moreover, this result holds independently of the number of (universal) spacetime dimensions and the VEV profile of the PQ field in the bulk. We, therefore, expect that more complex constructions, with additional bulk fields or less generic compactifications, are required to produce the most exotic patterns in Fig. 1. In turn, the observation of a plateau of heavy fields together with a canonical QCD axion could be explained by a single PQ field, propagating in the bulk of a hidden dimension. We will demonstrate how a potential observation of such a plateau could in this case hint towards the location of the zero mode in the canonical QCD axion band.
@@ II Extra dimensional models
with $y_{a}$ denoting a coordinate of the extra dimensions, that we assume to be compactified on a $(S^{1}/\mathbb{Z}^{2})$ orbifold of universal radius $R$. Accordingly, $y_{a}\to-y_{a}$ are identified for all points in the interval $[-\pi R,\pi R]$. In the equation above, $g$ is the determinant of the metric, $\mathcal{R}$ the Ricci scalar, $M_{d}$ the $d$-dimensional Planck mass, and $\Lambda_{B}$ the vacuum energy of the bulk  [^1] We will denote coordinates belonging to the full extra-dimensional models, to the 4D EFT and to the extra-dimensions with capital indices (e.g. $x^{A}$), greek letters (e.g. $x^{\mu}$ with $\mu=0,1,2,3$) and in latin lower case (e.g. $x^{a}$ with $a=5,6,\dots,\,d$, or $y_{a}$ with $a=1,\dots\,,\delta$), respectively. We will also make use of $\mathbf{y}=(y_{1},\dots\,,y_{\delta})$ to highlight the full set of extra-dimensional coordinates.. Regarding the matter content, we focus on models where the SM fields are localized on the branes, that is the boundaries of the extra dimensions.
@@
where the $\delta$-function localizes the axion couplings to QCD on the SM brane, at $y_{a}=\pi R$. The interactions above are assumed to be generated at a fundamental scale $M_{s}$, with $f_{d}$ denoting the PQ breaking scale. The former can be removed from the kinetic term by an appropriate redefinition of the axion field and the axion decay constant.
@@ III General mass matrix
$m_{\rm PQ}^{2}\equiv\chi_{\rm QCD}/f_{\rm PQ}^{2}$ is the standard PQ mass and
$\chi_{\text{QCD}}\equiv m_{\pi}^{2}f_{\pi}^{2}\frac{m_{u}m_{d}}{\left(m_{u}+m_{d}\right)^{2}}$ (III.3)
is the QCD topological susceptibility. Such mass matrix reproduces any of the models described in the previous section, by simply matching the WFs to Eqs. (II.17) (flat) and (II.28) (RS). Note that the main difference between these constructions in the $\mu\gg 1$ limit lies not in the mass ratios, but rather in the expressions for these WFs, which are exponentially enhanced by the curvature.
@@
${\sum\limits_{n=0}^{\infty}\dfrac{(\psi_{n})^{2}}{\lambda^{2}-(\mu_{n}/\mu_{1})^{2}\mathsf{y}^{2}}=1\,;}$ (III.5)
see App. A for details. We note that while this formula has been derived in Ref. [8] for the flat scenario, here we prove that it holds also for RS models. A more general approach to obtain the eigenvalues will also be derived in Sec. VI.
@@ IV Maximal displacements of KK axions
The $g$-factors can now be found using Eqs (IV.4) and (III.5):
@@
Remarkably, using the definition above, one rediscovers the QCD axion sum rule in the extra-dimensional models:
@@
proven in all generality in Ref. [5]. In the flat scenario, this result has been previously derived in Ref. [8]. In App. B we present a different proof, that extends beyond the flat case. We will now infer the consequences of these general results to the models introduced in the previous section.
@@ IV.1 QCD maxions in flat space
In flat scenarios, the sum in Eq. (III.5) can be performed exactly:
@@
These results confirm our expectations and agree with previous studies of the mass spectrum of this theory [18].
@@ IV.2 QCD maxions in curved space
In the RS model, the computations are more elaborate as $\mu_{n}\propto\gamma_{n}$. The sum in Eq. (III.5) can be performed exactly by means of the identity [25]
@@
These results, together with Eq. (IV.33), allow us to estimate the localization of the lightest eigenvalue, as well as of the plateau, for the several scenarios represented in Fig. 2(b). We note that a partial study of the mass spectrum of this model was previously presented in the literature [26], while expressions for the masses and some mixing factors were derived in Ref. [27].
@@ V Theoretical and experimental constraints
Figure 3: Collection of bounds on massive gravitons as a function of the lightest KK graviton mass, $\mu_{1}$. The red, green and purple regions are excluded by fifth force experiments [30, 31, 32, 33, 34, 35, 36, 37], astrophysics [38, 39] and collider searches [28, 40, 41], respectively. The black dashed line corresponds to the threshold mass value below which KK maxions can be generated; see Eq. (V.8). This assumes that the KK tower remains perturbative up to TeV energies.
To combine these constraints, motivated by the form of the eigenvalue equation (III.5), we consider a simplified approach to extra-dimensional maxions, that relies on using the $2\times 2$ block mass matrix
@@
This is equivalent to truncating the sum in Eq. (III.5) to the second term. The lightest eigenvalue can then be predicted up to an accuracy of $\mathcal{O}(\mu_{1}/\mu_{2})^{2}$ and, in the limit $\mathsf{y},\psi_{0}\ll\psi_{1}$, reads:
@@
we obtain $g_{a\gamma\gamma\,,\text{plateau}}\approx\,{2.5\times 10^{-10}}\,{\text{GeV}}^{-1}$, after fixing $\mu_{1}\sim 10^{-2}$ eV, $\Lambda\sim 1$ TeV and $|C_{a\gamma}|\approx 1.92$ since all axions inherit the same model-independent coupling to photons from the original interaction in Eq. (II.4). Such coupling is on the verge of the constraints imposed by single-axion searches in stars; see Fig. (1). We note that model-dependent interactions could be introduced in the IR brane to make the KK axions more photophobic, at the cost of some tuning (see e.g. Ref. [42]). Nevertheless, supernova data probes directly the gluonic coupling in Eq. (V.10) for the largest possible displacements identified in Eq. (V.9). Such bound takes into account the production of a single axion in the star, which does not hold in our setup. However, as we show below, the production of additional eigenstates strengthens this limit severely in the parameter space of interest.
Figure 4: The “rescaling factor” $\aleph(E_{c})$, defined in Eq. (V.13), as a function of the PQ mass. The different lines correspond to different benchmark values for $\mu_{1}$ and $\psi_{1}$ that allow for maxion regimes interpolating the RS and flat scenarios. The characteristic energy was taken to be $E_{c}=30$ MeV. The parameter space to the left of the dashed lines, where $\Lambda>1$ TeV, is consistent with the EFT analysis; see the text for details.
By employing the approach of Ref. [20], one can obtain the effective interaction scale of the KK axions, integrated up to some characteristic energy $E_{c}$ (around $30$ MeV for typical supernovae core temperatures):
$\frac{1}{f_{4,\text{eff.}}}\equiv\frac{\aleph(E_{c})}{f_{4}}\,,$ (V.12)
where
$\aleph^{2}(E_{c})\equiv f_{4}^{2}\sum\limits_{(\lambda m_{4})<E_{c}}\frac{1}{f_{\lambda}^{2}}=\sum\limits_{(\lambda m_{4})<E_{c}}\left(\frac{\lambda^{2}}{g_{\lambda}}\right)\,.$ (V.13)
Let us now take some interesting limits of this expression, assuming $E_{c}\gg\mu_{1}$. In flat scenarios, we obtain:
$\aleph_{\text{flat}}^{2}(E_{c})\approx 2\times\begin{cases}\frac{1}{3\pi^{2}}\left(\frac{\mu_{1}E_{c}^{3}}{m_{4}^{4}}\right)&m_{4}\gg\mu_{1}\,,\\ \left(\frac{E_{c}}{\mu_{1}}\right)&m_{4}\ll\mu_{1}\,;\end{cases}$ (V.14)
while in the warped case,
$\aleph_{\text{RS}}^{2}(E_{c})\approx\begin{cases}\frac{4}{3\pi\gamma_{1}}\left(\frac{\mu_{1}E_{c}^{3}}{\psi_{1}^{2}m_{4}^{4}}\right)&\psi_{1}m_{4}\gg\mu_{1}\,,\\ \frac{\gamma_{1}}{\pi}\left(\frac{\psi_{1}^{2}\,E_{c}}{\mu_{1}}\right)&\psi_{1}m_{4}\ll\mu_{1}\,.\end{cases}$ (V.15)
The results in Eq. (V.14), applicable to flat extra-dimensional models, agree with previous findings reported in Ref. [20]. In maxion-dominated regimes, $\aleph_{\rm flat}^{2}\sim\mathsf{y}\lesssim 1$, so that the collective KK bounds can become weaker relative to those obtained for a single axion. In contrast, in the regime dominated by the plateau modes, the relation $1/f_{4,\text{eff.}}\sim\sqrt{N}/f_{4}$, where $N\approx E_{c}/\mu_{1}$, leads to a substantial strengthening of these constraints. Note that the transition between this mass-dependent and flat regimes occurs approximately at
$m_{4}^{\star}\approx\frac{\sqrt{\mu_{1}E_{c}}}{2\,\psi_{1}}\,.$ (V.16)
By extending these results to the warped case, we find that the curvature exponentially suppresses the maxion regimes, while it enhances the plateau contribution. Whichever of these effects dominates, and therefore the fate of the bounds, depends on the parameter space of interest.
Figure 4 shows the values of the “rescaling factor”, $\aleph$, computed numerically for the flat and different RS scenarios. In this figure, we can clearly identify the transition between maxion and plateau dominated regimes, as well as other features discussed previously. Choosing $\mu_{1}=10^{-2}$ eV, which produces the largest $g$-factors allowed by fifth force constraints in the flat scenario (see Eq. (V.9)), we find that the rescaling factor is dominated by the plateau modes and, therefore, it is extremely large, $\aleph_{\rm flat}\sim 10^{5}$. To compare this with curved models, we have chosen a larger value for $\mu_{1}=1$ eV, that complies with gravity constraints for $\psi_{1}$ as large as $10^{4}$. For the same $\Lambda=1$ TeV (represented by the vertical lines in Fig. 4), we infer that the warped scenario worsens significantly the bounds, with $\aleph_{\rm RS}\sim\aleph_{\rm flat}\psi_{1}\sim 10^{8}$. Such a benchmark corresponds already to a very small displacement of the KK modes from the QCD canonical line, $g_{0}\sim 10^{2}$.
We can take the argument further by computing the minimum value of $m_{4}$, in Eq. (V.15), which is compatible with a suppression of the bounds i.e. $\aleph(E_{c})<1$. Plugging in this value in the expression for the maxion $g$-factor, we obtain:
$m_{4}^{2}\gtrsim\left(\frac{\mu_{1}E_{c}^{3}}{\psi_{1}^{2}}\right)^{1/2}$$\implies$ $g_{0}\approx\frac{\psi_{1}^{2}}{\mathsf{y}^{2}}\gtrsim\psi_{1}\left(\frac{E_{c}}{\mu_{1}}\right)^{3/2}\,.$ (V.17)
Written in terms of the cutoff scale, as in Eq. (V.7), such condition implies that
$\mu_{1}$ $\gtrsim\frac{\psi_{1}^{2}}{(36\pi)^{6}}\frac{\Lambda^{6}E_{c}^{3}}{\chi_{\text{QCD}}^{2}}\,$ (V.18)
$\approx 10^{10}\,\psi_{1}^{2}\,\left(\frac{\Lambda}{1\leavevmode\nobreak\ \text{TeV}}\right)^{6}\left(\frac{E_{c}}{30\leavevmode\nobreak\ \text{MeV}}\right)^{3}\leavevmode\nobreak\ \text{GeV}\,,$
which is clearly incompatible with the maxion regime. Indeed, the equation above is in strong contradiction with Eq. (V.8). This shows unambiguously that $\aleph(E_{c})>1$ even for maxion regimes, unless $E_{c}$ is comparable to $\mu_{1}$.
@@
We, therefore, conclude that the maxion patterns from the extra-dimensional models analysed in this work are in severe tension with the joint combination of gravity and astrophysical constraints.